OpenStudy (anonymous):

Simplify: the quantity m over 3, this quantity over the quantity 8 times m squared all over 15, end quantity. 5 over the quantity 8 times m squared, end quantity., m ≠ 0 5 over the quantity 8 times m, end quantity., m ≠ 0 8 times m squared all over 5., m ≠ 0 8 times m all over 5., m ≠ 0

OpenStudy (anonymous):

@FutureMathProfessor heres the prob

OpenStudy (anonymous):

@xoya please help

OpenStudy (anonymous):

\[\left(\frac{m}{3}\right)/\left(\frac{8 m^2}{15}\right)=\frac{5}{8 m} \]